Please help me solve this:
y = 2x - 2
3x + 4y = 3
thank you.

Found 2 solutions by mananth, BGutridge:
Answer by mananth(16946)

-2 x + 1 y = -2 .............1

3 x + 4 y = 3 .............2
Eliminate y
multiply (1)by -4
Multiply (2) by 1
8 x -4 y = 8
3 x + 4 y = 3
Add the two equations
11 x = 11
/ 11
x = 1
plug value of x in (1)
-2 x + 1 y = -2
-2 + y = -2
y = -2 + 2
y = 0
y = 0
(1,0) is the solution

Answer by BGutridge(15)

Since one equation is already solved for "y", plug that expression in for "y" in the other equation 3x+4(2x-2)=3. Then simplify, leaving:
3x+8x-8=3
11x=11
x=1
Plugging x back into the other equation, you get y=2(1)-2, or y=0, so the complete solution is (1,0)